OpenStudy (stormoshawty):

Make the equation in slope-intercept form. -0.6x - 0.8y = 0.30

OpenStudy (the_fizicx99):

:P -0.6x - 0.8y = 0.30 +.6x +.6x -.8y = 0.6x + 0.30 /-.8 /-.8 /-.8 y = -.6/.8x -0.375

OpenStudy (stormoshawty):

I think you messed up somewhere in there .-.

OpenStudy (kc_kennylau):

The approach is simply to make \(y\) the subject :)

OpenStudy (the_fizicx99):

I didn't. I showed the process. You need to simplify -.6/.8x

OpenStudy (the_fizicx99):

divide -.6/.8

OpenStudy (the_fizicx99):

What did you get?

OpenStudy (stormoshawty):

0.75

OpenStudy (the_fizicx99):

negative .6/.8 = -.75 :D So whats your y = mx+b now

OpenStudy (stormoshawty):

y=-0.75x-0.375

OpenStudy (the_fizicx99):

Yup :)